There are 7 girls and 6 boys in a class. Suppose five children are chosen randomly to decorate the room. What is the probability that exactly two boys are chosen?
Answer by ewatrrr(24785)


Hi,
7 girls and 6 boys in a class, five children are chosen
p(exactly 2 boys) = 6C2 * 7C3 / 13C5 = 15*35/1287 = .4079